Please Solve Showing all steps.

CT Use Romberg Integration to evaluate S cos x dx Find R5,5 (Note true value is 0.08871 )
Answer #1

J coszor Ri,, - . [cos o + cos n] = 0 R2,1 = at [Cos 0 + 2005! + cos n] = 0 Rg, 1 - [ coso +2lcos+ cos+ 605 3.) + costs] ; l.: R4,4 = R4,3 +63(R 4,3 -R3,5) = 0 R5,4 - 8,3 114 ( ,5 - 24,5) = 0 .: R$5 = R5A + L (R5; 4 -R44) = 0 Rs.=0 CScanned with Cam